# Current affiliation:] Laboratoire Charles Coulomb (L2C), Univ. Montpellier, CNRS, Montpellier, France

### v1

@@ Abstract
where $\mathcal{R}\equiv 4V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}+(V_{\textsl{{a}}\textsl{{a}}}-V_{\textsl{{b}}\textsl{{b}}})^{2}$. When $\mathcal{R}>0$, the eigenvalues in Eq. (5) are purely imaginary, and the stability of the solution to long-wavelength perturbations is determined at ${\cal O}(|\bm{q}|^{2})$. On the other hand, for $\mathcal{R}<0$ —i.e., when $V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}<-\tfrac{1}{4}(V_{\textsl{{a}}\textsl{{a}}}-V_{\textsl{{b}}\textsl{{b}}})^{2}$— the eigenvalue $\Lambda_{-}$ acquires a positive real part, indicating the emergence of an instability. In particular, for reciprocally aligning populations $V_{\textsl{{a}}\textsl{{b}}}$ and $V_{\textsl{{b}}\textsl{{a}}}$ share the same sign and ${\cal R}>0$, while the condition $V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}<0$ requires a sufficient amount of nonreciprocity. Anti-flocking mixtures are more difficult to analyze, as in this case the instability is not purely longitudinal or transverse [Fig. 3(b)]. Nevertheless, the solid pink lines in Figs. 3 enclosing regions for which $\mathbb{M}_{\rm h}$ has an eigenvalue with positive real part show excellent agreement with the direct numerical estimate based on the six-dimensional system. In particular, for $\bar{\chi}>0$ the pink line corresponds to ${\cal R}=0$ when the instability is longitudinal.
@@
We have shown that even weakly nonreciprocal interactions significantly impact the collective dynamics of flocking active matter. Combining agent-based simulations and the analysis of a coarse-grained theory, our results demonstrate that nonreciprocal alignment alone constitutes a generic route to demixing, without resorting to other mechanisms such as population heterogeneity [14] or pairwise repulsion [29, 37]. Our theoretical study further emphasises that this phenomenology is induced by the coupling between density and orientational order, which shapes most of the characteristics of flocks. Our findings echo recent works highlighting the fragility of flocks [58] and their sensitivity to the presence of spatial anisotropy [59], inclusions and boundaries [60, 61], and quench disorder [62, 63, 64]. Since they rely solely on weakly nonreciprocal alignment, we believe that our results will have implications for the modeling of a variety of systems, including mixtures of Quincke rollers [14] or Janus colloids [30], and migrating heterogeneous cellular tissues [35].

### v2

@@ Abstract
where $\mathcal{R}\equiv 4V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}+(V_{\textsl{{a}}\textsl{{a}}}-V_{\textsl{{b}}\textsl{{b}}})^{2}$. When $\mathcal{R}>0$, the eigenvalues in Eq. (5) are purely imaginary, and the stability of the solution to long-wavelength perturbations is determined at ${\cal O}(|\bm{q}|^{2})$. On the other hand, for $\mathcal{R}<0$ —i.e., when $V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}<-\tfrac{1}{4}(V_{\textsl{{a}}\textsl{{a}}}-V_{\textsl{{b}}\textsl{{b}}})^{2}$— the eigenvalue $\Lambda_{-}$ acquires a positive real part, indicating the emergence of an instability. For reciprocally aligning populations $V_{\textsl{{a}}\textsl{{b}}}$ and $V_{\textsl{{b}}\textsl{{a}}}$ share the same sign and ${\cal R}>0$, while the condition $V_{\textsl{{a}}\textsl{{b}}}V_{\textsl{{b}}\textsl{{a}}}<0$ requires a sufficient amount of nonreciprocity. Anti-flocking mixtures are more difficult to analyze, as in this case the instability is not purely longitudinal or transverse [Fig. 3(b)]. Nevertheless, the solid pink lines in Figs. 3 enclosing regions for which $\mathbb{M}_{\rm h}$ has an eigenvalue with positive real part show excellent agreement with the direct numerical estimate based on the six-dimensional system. When the instability is longitudinal, these lines correspond to ${\cal R}=0$.
@@
We have shown that even a weak amount of nonreciprocity significantly impacts the collective dynamics of flocking active matter. Combining agent-based simulations and the analysis of a coarse-grained theory, our results demonstrate that nonreciprocal alignment alone constitutes a generic route to demixing, without resorting to other mechanisms such as population heterogeneity [39] or pairwise repulsion [32, 31]. Our theoretical analysis further emphasises that this phenomenology is induced by the coupling between density and orientational order, which shapes most of the characteristics of flocks. Our findings echo recent works highlighting the metastability of flocks [17, 5] and their fragility to the presence of spatial anisotropy [50], inclusions and boundaries [17, 24], quench [44, 11, 22] and chirality [58] disorder. Since they rely solely on weakly nonreciprocal alignment, we believe that our results will have implications for the modeling of a variety of systems, including mixtures of Quincke rollers [39] or Janus colloids [57], and migrating heterogeneous cellular tissues [49].